# Regularization by regular noise: a numerical result

### v1

@@ Abstract.
We study a singular stochastic equation driven by a regular noise of fractional Brownian type with Hurst index $H\in(1,\infty)\setminus\mathbb{Z}$ and drift coefficient $b\in\mathcal{C}^{\alpha}$, where $\alpha>1-\tfrac{1}{2H}$. The strong well-posedness of this equation was first established in [Ger23], a phenomenon referred to as regularization by regular noise. In this note, we provide a corresponding numerical analysis. Specifically, we show that the Euler–Maruyama approximation $X^{n}$ converges strongly to the unique solution $X$ with rate $n^{-1}$. Furthermore, under the additional assumption $b\in\mathcal{C}^{1}$, we show that $n(X-X^{n})$ converges to a non-trivial limit as $n\to\infty$, thereby confirming that the rate $n^{-1}$ is in fact optimal upper bound for this scheme.
@@ 1. Introduction
It is known that when $H\in(0,1)$, a $d$-dimensional fractional Brwonian motion (fBM) $B^{H}$ can be defined via the Mandelbrot - van Ness representation [MVN68]:
@@
Due to the possible multiple times of integral in (1.2) we can see clearly that the path of $B^{H}$ for $H>1$ is regular, at least $C^{1}$. This fact yields one of the appealing points of [Ger23] which shows the strong well-posedness of following equation with singular $b$:
@@
which shares the similar framework from [CdRHM22, HZZZ24, HLL24, Lin24] but with non-Markvin noise when $H\neq k+\frac{1}{2}$ for any $k\in{\mathbb{N}}$.
In this note our motivation is to provide a numerical result of it by considering its Euler-Maruyama (EM) scheme
@@
When $H\in(0,1)$, [BDG21] has shown the strong convergence rate $n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}$ of the scheme (1.9) to (1.3), inside which the Girsanov Theroem and Stochastic Sewing Lemma (SSL) [Lê20] play one of the most important roles. Precisely the methodology inside [BDG21] can be roughly summarized as (take $x_{0}=x_{0}^{n}$)
@@
When $H>1$, on one hand, as indicated in [Ger23, GG25] the Girsanov Theorem becomes less helpful; meanwhile PDE tools clearly would not work. Therefore the challenge for showing the convergence of (1.9) to (1.3) lies in two folds comparing with known results: absence of Girsanov and non-Markovianity. Alternatively, although [Ger23] studies the well-posedness only, actually it already hints the possible way to get rid of Girsanov; concerning numerical approximation [BDG24] provides the similar evidence, however it considers the singular SDE driven by an $\alpha$ stable process which is a Markov process. Nevertheless we are able to show convergence rate $n^{-1}$ which is comparable with [BDG21] since $H>1$ here implies $\alpha>1-\frac{1}{2H}>\frac{1}{2}$.
For such obtained convergence rate, naturally we wonder how far it is away from being optimal. Such question on the optimality of EM scheme has been addressed and studied for instance in [MGY23, EMGY25, DGL23a] concerning Brownian noise, [PSS24] for Lévy processes with jumps and [Neu06, NN07, LLT25, HLN16] for fBM with $H<1$. In particular [Neu06] confirms for $H\in(\frac{1}{2},1)$, order $n^{H}$ reaches its optimality of EM scheme (1.9) for $b\in{\mathcal{C}}^{2}$ (twice differentiable) with possible linear growth. Here we complete this result with rate $n^{-1}$ for $H>1$ and less regular $b$ (actually only ${\mathcal{C}}^{1}$). The idea of the proof is straightforward: we show that for $b\in{\mathcal{C}}^{1}$, the following approximation reaches a possibly non-zero limit:
$n(X_{t}-X^{n}_{t})\overset{n\rightarrow\infty}{\rightarrow}c(t)\neq 0,\quad a.s.$
which indicates that for (1.9) the best convergence rate we can expect should be not faster than order $n$. Evidently it also verifies that the rate we obtained actually is optimal.
Lastly, let us mention a few related works that share a similar interest and spirit in the study of numerical approximations. For equations of the type (1.3) with singular drift, convergence results have been established in [NS21, DG20, JM21, BW25] for additive Brownian motion, in [DGL23b, BDG21, GLL25, LL25] for multiplicative Brownian noise, in [BDG24, BWWZ25] for Lévy processes with jumps, and in [BDG21, GHR25] for fractional Brownian motion. A slightly different notion of singularity-namely, piecewise Lipschitz coefficients-has also been investigated, with convergence results obtained in works such as [LS16, MGY20, MGRY25, PSS24]. We must acknowledge that this is only a brief selection from a vast body of literature, and we encourage readers to explore the cited works for further details and insights.
@@
In Section˜2 we introduce the necessary notations and main results. Section˜3 collects all of the crucial tools and properties of fBM. We present the central analysis and proof of the main result in Section˜4. In the end we discuss optimality in Section˜5.
@@ 2.1. Preliminaries
For $k\in\mathbb{N}$, $f:\mathbb{R}^{d}\mapsto\mathbb{R}$, denote $\partial_{k}f(x)\coloneqq\frac{\partial f(x)}{\partial x_{k}}$ for $x\in\mathbb{R}^{d}$ and $\nabla f(x)\coloneqq(\partial_{i}f(x))_{1\leqslant i\leqslant d}$, the derivative is understood in the weak sense. For vector-valued $f$ we use the same notation, and $\nabla^{k}f$ is defined via $\nabla(\nabla^{k-1}f)$ iteratively. For a multi-index $k=(k_{1},\ldots,k_{d})\in{\mathbb{N}}^{d}$, denote $\partial^{k}f(x)\coloneqq\frac{\partial^{|k|}f(x)}{\partial x_{k_{1}}\cdots\partial x_{k_{d}}}$. If $k=(0,\ldots,0)$, we use convention $\partial^{k}f=f$. We denote by ${C}^{\infty}_{0}$ (${C}_{p}^{\infty}$, resp.) the set of all continuously infinitely differentiable functions that, along with all of their partial derivatives, are compactly supported (of polynomial growth, resp.).
For $\alpha\in(0,1]$, we set $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ to be the space of continuous functions such that
@@ 2.2. Main results
Our main assumption and result can be stated as follows.
@@
Notice that following [Ger23], under the assumption above, there exists a unique strong solution to (1.3). Here is our numerical approximation result for it.
@@
Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly. Suppose Assumption˜2.1 holds. Then for every $p>1$, we have
$\|X-X^{n}\|_{C_{p}^{\frac{1}{2}}[0,1]}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1},$ (2.4)
where $N=N(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$\big\|\sup_{t\in[0,1]}|X_{t}-X_{t}^{n}|\big\|_{L_{\omega}^{p}}\leqslant N|x_{0}-x_{0}^{n}|+Nn^{-1}.$
Idea of the analysis. Its detailed proof will be given in Section˜4. Here we only outline the rough idea.
Observe that for any $p\geqslant 1$, for $0\leqslant s\leqslant t\leqslant 1$
$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})|\big\|_{L^{p}_{\omega}}=\big\|(\varphi_{t}-\varphi_{t}^{n})-(\varphi_{s}-\varphi_{s}^{n})\big\|_{L^{p}_{\omega}},$
@@
meanwhile, we know that our aim is to get rid of ${\mathcal{C}}^{1}$-regularity demand on $b$ via proper implement of the regularization from $B^{H}$, precisely the Gaussian density $p_{c(H)t^{2H}}$ which is infinitely many smoothing (see Section˜3.2). Also keep in mind in the current setting Girsanov Theorem is not available. What we have learned from [Ger23, BDG24] is that we can achieve this goal via freezing the exponent $\varphi_{r}$ inside the integral $\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ (the same for $\varphi^{n}_{k_{n}(r)}$ inside the integral $\int_{0}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$) via taking conditional expectation in the framework of SSL. That is to say, heuristically for tst-s being small enough
@@
and this $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}$” is verified by SSL (see Lemma˜3.1 below) by taking
@@
similarly for $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$, then together with the property of Gaussian density $p_{c(H)t^{2H}}$ of fBM ( $\mathcal{P}_{t}^{H}f:=p_{c(H)t^{2H}}\ast f$) we can further write
@@
Now we can see that instead of dealing with $b$ directly, we have more regularity to spare in $\mathcal{P}^{H}_{t}b$ due to the smoothing effect from the convolution with $p_{c(H)t^{2H}}$.
Although the true analysis later also contains large part technical settlement, the core of the whole argument is clear. In the end we are able to adjust the idea above into the proof of the convergence rate in the following form:
$\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|\varphi_{t}-\varphi_{t}^{n}\big\|_{L^{p}_{\omega}}$ $\leqslant N(\|\varphi-\varphi^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
$=N(\|X-X^{n}\|_{C^{\frac{1}{2}}_{p}}+n^{-1})|t-s|^{\frac{1}{2}+\varepsilon}$
for sufficiently small $\varepsilon$. Therefore we get (2.4) after fine dividing the interval $[0,1]$ and for each sub-interval applying what we have obtained above.

### v2

@@ Abstract.
We study a singular stochastic equation driven by regular noise of fractional Brownian type with Hurst index $H\in(1,\infty)\setminus\mathbb{Z}$ and drift coefficient $b\in\mathcal{C}^{\alpha}$, where $\alpha>1-\tfrac{1}{2H}$. The strong well-posedness of this equation was first established in [13], a phenomenon known as regularization by regular noise. In this note, we provide a numerical analysis of the equation. Specifically, we prove that the Euler–Maruyama approximation $X^{n}$ converges strongly to the unique solution $X$ at rate $n^{-1}$. Moreover, we show that $n(X-X^{n})$ converges in probability to a non-trivial limit as $n\to\infty$, which confirms that the rate $n^{-1}$ is optimal for this scheme. In this sense, this provides a first-order numerical method for equations with non-Lipschitz drift while still achieving the rate $n^{-1}$.
@@ 1. Introduction
It is known that when $H\in(0,1)$, a $d$-dimensional fractional Brownian motion (fBM) $B^{H}$ can be defined via the Mandelbrot–van Ness representation [26]:
@@
Due to the possible multiple integrals in (1.2), we can see clearly that the paths of $B^{H}$ for $H>1$ are regular, at least $C^{1}$. This fact leads to one of the appealing results of [13], which shows the strong well-posedness of the following equation with singular $b$:
@@
which shares a similar framework with [6, 16, 15, 25], but with non-Markovian noise when $H\neq k+\frac{1}{2}$ for any $k\in{\mathbb{N}}$.
Our motivation here is to provide a numerical result for this equation by considering its Euler–Maruyama (EM) scheme
@@
When $H\in(0,1)$, [4] has shown the strong convergence rate $n^{-(\frac{1}{2}+\alpha H)\wedge 1+\epsilon}$ of the scheme (1.9) to (1.3), in which the Girsanov Theorem and the Stochastic Sewing Lemma (SSL) [22] play crucial roles. More precisely, the methodology in [4] can be roughly summarized as follows (taking $x_{0}=x_{0}^{n}$):
@@
When $H>1$, on the one hand, as indicated in [13, 11], the Girsanov Theorem becomes less helpful; meanwhile, PDE tools clearly do not apply. Therefore, the challenge in showing the convergence of (1.9) to (1.3) lies in two aspects compared with known results: the absence of Girsanov’s theorem and the non-Markovian nature of the noise. Alternatively, although [13] studies only well-posedness, it already hints at a possible way to avoid using Girsanov’s theorem; concerning numerical approximation, [3] provides similar evidence, but it considers singular SDEs driven by an $\alpha$-stable process, which is Markovian. Nevertheless, we are able to show a convergence rate of $n^{-1}$, which is comparable with [4], since $H>1$ here implies $\alpha>1-\frac{1}{2H}>\frac{1}{2}$.
Given this convergence rate, it is natural to ask how far it is from being optimal. Such questions on the optimality of the EM scheme have been addressed, for instance, in [31, 10, 7] for Brownian noise, [35] for Lévy processes with jumps, and [34, 32, 24, 19] for fBM with $H<1$. In particular, [34] confirms that for $H\in(\frac{1}{2},1)$, the order $n^{H}$ is optimal for the EM scheme (1.9) when $b\in{\mathcal{C}}^{2}$ (twice differentiable) with possible linear growth. Here, we complete this result with rate $n^{-1}$ for $H>1$ and less regular $b$ (in fact, only ${\mathcal{C}}^{1}$). The idea of the proof is straightforward: we show that for $b\in{\mathcal{C}}^{\alpha}$, the following approximation converges in probability to a possibly non-zero limit:
$n(X_{t}-X^{n}_{t})\overset{n\rightarrow\infty}{\rightarrow}c(t)\neq 0,$
which indicates that for (1.9), the best convergence rate one can expect is no faster than order $n^{-1}$. Consequently, this also verifies that the rate we obtain is indeed optimal. Evidently, (1.9) provides a stochastic numerical method for simulating equations with singular $b$ while still converging with rate $n^{-1}$.
Lastly, we mention a few related works that share a similar interest and spirit in the study of numerical approximations. For equations of the type (1.3) with singular drift, convergence results have been established in [33, 9, 20, 2] for additive Brownian motion, in [8, 4, 12, 21] for multiplicative Brownian noise, in [3, 1] for Lévy processes with jumps, and in [4, 14] for fractional Brownian motion. A slightly different notion of singularity-namely, piecewise Lipschitz coefficients-has also been investigated, with convergence results obtained in works such as [23, 30, 29, 35]. We emphasize that this is only a brief selection from a vast body of literature, and we encourage readers to consult the cited works for further details and insights.
@@
In Section˜2 we introduce the necessary notations and main results. Section˜3 collects all of the crucial tools and properties of fBM. We present the central analysis and the proof of the main convergence result in Section˜4. Finally, we show optimality in Section˜5. Appendix˜A contains several auxiliary proofs.
@@ 2.1. Preliminaries
For $k\in\mathbb{N}$, $f:\mathbb{R}^{d}\mapsto\mathbb{R}$, denote $\partial_{k}f(x)\coloneqq\frac{\partial f(x)}{\partial x_{k}}$ for $x\in\mathbb{R}^{d}$ and $\nabla f(x)\coloneqq(\partial_{i}f(x))_{1\leqslant i\leqslant d}$, the derivative is understood in the weak sense. For vector-valued $f$ we use the same notation, and $\nabla^{k}f$ is defined via $\nabla(\nabla^{k-1}f)$ iteratively. For a multi-index $k=(k_{1},\ldots,k_{d})\in{\mathbb{N}}^{d}$, denote $\partial^{k}f(x)\coloneqq\frac{\partial^{|k|}f(x)}{\partial x_{k_{1}}\cdots\partial x_{k_{d}}}$. If $k=(0,\ldots,0)$, we use convention $\partial^{k}f=f$. We denote by ${{\mathcal{C}}}^{\infty}_{0}$ (${{\mathcal{C}}}_{p}^{\infty}$, resp.) the set of all continuously infinitely differentiable functions that, along with all of their partial derivatives, are compactly supported (of polynomial growth, resp.).
For $\alpha\in(0,1)$, we set $\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ to be the space of continuous functions such that
@@
Let $n\in{\mathbb{N}}$ and $\alpha,\beta\in(0,1)$ such that $\alpha+\beta>1$. Then for $f\in{\mathcal{C}}^{\alpha}([0,T],{\mathbb{R}}^{n})$, $g\in{\mathcal{C}}^{\beta}([0,T],{\mathbb{R}}^{n})$, the Young integral $h_{t}=\int_{0}^{t}f_{t}dg_{t}$ is well-defined as the limit as $m\rightarrow\infty$ of the Riemann sums
$\sum_{i=0}^{m}f_{t_{i}^{m}}(g_{t_{i+1}^{m}\wedge t}-g_{t_{i}^{m}\wedge t})$
where $(t_{i}^{m})_{0\leqslant m}$ is any partition sequence of $[0,t]$. And the Young integral satisfies the estimates: there exists the constant $C$ depends only on $\alpha,\beta$ so that for all $0\leqslant s\leqslant t\leqslant T,$
$|h_{t}-h_{s}-f_{s}(g_{t}-g_{s})|\leqslant C|t-s|^{\alpha+\beta}[f]_{{\mathcal{C}}^{\alpha}([s,t])}[g]_{{\mathcal{C}}^{\beta}([s,t])},$
which also yields the following
$[h]_{{\mathcal{C}}^{\beta}([s,t])}\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}([s,t])}[g]_{{\mathcal{C}}^{\beta}([s,t])}.$ (2.1)
@@ 2.2. Main results
By scaling we can always take $t\in[0,1]$ without lost of generality. Our main assumption and results are stated as follows.
@@
Notice that following [13], under the assumption above, there exists a unique strong solution to (1.3). Here is our numerical approximation result for it.
@@
Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly. Suppose Assumption˜2.1 holds. Then for every $p\geqslant 1$ and $\gamma<1+(\alpha-1)H$, we have
$\|X-X^{n}\|_{C_{p}^{\gamma}[0,1]}\leqslant C|x_{0}-x_{0}^{n}|+Cn^{-1},$ (2.5)
where $C=C(p,d,\alpha,H,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$\big\|\sup_{t\in[0,1]}|X_{t}-X_{t}^{n}|\big\|_{L_{\omega}^{p}}\leqslant C|x_{0}-x_{0}^{n}|+Cn^{-1}.$
The following theorem settles the question on optimality.
Theorem 2.4.
Let $(X_{t})_{t\in[0,1]},(X_{t}^{n})_{t\in[0,1]}$ be the solutions to (1.3) and (1.9) accordingly and $x_{0}=x_{0}^{n}$. Suppose Assumption˜2.1 holds. Then for every $f\in{\mathcal{C}}^{\alpha}({\mathbb{R}}^{d})$ there exists a $({\mathcal{F}}_{t})_{t\in[0,1]}$-adapted process $({\mathcal{H}}f)^{X}_{t\in[0,1]}\in C_{p}^{1+(\alpha-1)H-}[0,1]$ for every $p\geqslant 2$ such that bounded linearly
${\mathcal{C}}^{\alpha}\ni f\mapsto({\mathcal{H}}f)^{X}\in C_{p}^{1+(\alpha-1)H-}[0,1]$
and if $g\in{\mathcal{C}}^{1}({\mathbb{R}}^{d})$, with probability one
$({\mathcal{H}}g)_{t}^{X}=\int_{0}^{t}\nabla g(X_{s})\mathrm{d}s,\quad t\in[0,1];$
moreover, we have for any $t\in[0,1]$, in probability
$\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})=:c(t)$ (2.6)
exists and $c\in{\mathcal{C}}^{1+(\alpha-1)H-}([0,1])$ satisfies
$c(t)=\int_{0}^{t}c(s)\mathrm{d}({\mathcal{H}}b)^{X}_{s}+\frac{1}{2}(b(X_{t})-b(x_{0})),\quad t\in[0,1],$ (2.7)
where the integral inside (2.7) is defined in the sense of Young.
Idea of the analysis. The detailed proofs of Theorem˜2.2 and Theorem˜2.4 will be given in Section˜4 and Section˜5, respectively. Here we only outline the main ideas.
  1. (1)
    (Convergence) Observe that for any $p\geqslant 1$, for $0\leqslant s\leqslant t\leqslant 1$,
    $\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|(\varphi_{t}-\varphi_{t}^{n})-(\varphi_{s}-\varphi_{s}^{n})\big\|_{L^{p}_{\omega}},$
@@
    meanwhile, we know that our aim is to remove the ${\mathcal{C}}^{1}$ regularity requirement on $b$ via proper use of the regularization coming from $B^{H}$, namely the Gaussian density $p_{c(H)t^{2H}}$, which is infinitely smoothing (see Section˜3.2). Also keep in mind that in the current setting the Girsanov Theorem is not available. What we have learned from [13, 3] is that we can achieve this goal via freezing the exponent $\varphi_{r}$ inside the integral $\int_{0}^{t}b(\varphi_{r}+B_{r}^{H})\mathrm{d}r$ (and similarly for $\varphi^{n}_{k_{n}(r)}$ inside the integral $\int_{0}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$) by taking conditional expectation in the framework of the SSL. That is to say, heuristically, for tst-s small enough,
@@
    and this $\overset{\|\cdot\|_{L^{p}_{\omega}}}{\approx}$” is justified by the SSL (see Lemma˜3.1 below) by taking
@@
    and similarly for $\int_{s}^{t}b(\varphi^{n}_{k_{n}(r)}+B_{r}^{H})\mathrm{d}r$. Then, together with the property of the Gaussian density $p_{c(H)t^{2H}}$ of fBM ( $\mathcal{P}_{t}^{H}f:=p_{c(H)t^{2H}}\ast f$), we can further write
@@
    Now we can see that instead of dealing with $b$ directly, we gain additional regularity in $\mathcal{P}^{H}_{t}b$ due to the smoothing effect of the convolution with $p_{c(H)t^{2H}}$.
    Although the full analysis later also contains many technical details, the core of the argument is clear. In the end, we are able to turn the idea above into a proof of the convergence rate in the following form:
    $\big\|(X_{t}-X_{t}^{n})-(X_{s}-X_{s}^{n})\big\|_{L^{p}_{\omega}}=\big\|\varphi_{t}-\varphi_{t}^{n}\big\|_{L^{p}_{\omega}}$ $\leqslant C(\|\varphi-\varphi^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon}$
    $=C(\|X-X^{n}\|_{C^{\gamma}_{p}[s,t]}+n^{-1})|t-s|^{\gamma+\varepsilon}$
    for sufficiently small $\varepsilon$. Therefore, we obtain (2.5) after finely dividing the interval $[0,1]$ and applying the above estimate on each sub-interval.
  2. (2)
    (Optimality) Our idea for verifying optimality is to establish a limit theorem for the asymptotic error distribution of $X^{n}$ and $X$, that is, to show that $n(X^{n}-X)$ converges to a nontrivial limit, indicating that order $n$ is the optimal convergence rate of $X^{n}$ from (1.9) to $X$.
    For this, morally, if the coefficient $b$ is smooth, we can show that there exists a nonzero process $c(t)$ such that $c(t)=\lim_{n\rightarrow\infty}n(X_{t}-X_{t}^{n})$ in probability and it satisfies a random ODE
    $c(t)=\int_{0}^{t}\nabla b(X_{s})c(s)ds+\frac{1}{2}(b(X_{t})-b(x_{0})).$ (2.8)
    At this point, we already see the difficulty in our setting: $b\in{\mathcal{C}}^{\alpha}$ with $\alpha\in(1-\frac{1}{2H},1)$ implies that $\nabla b(X)$ is ill-defined in the above equation.
    However, if we reformulate (2.8) equivalently as follows: denote $H_{t}:=\int_{0}^{t}\nabla b(X_{s})\mathrm{d}s$,
    $c(t)=\int_{0}^{t}c(s)\mathrm{d}H_{s}+\frac{1}{2}(b(X_{t})-b(x_{0})),$ (2.9)
    then, due to the regularization effect, we can show that a.s. $H$ is ${\mathcal{C}}^{1+(\alpha-1)H-}$-Hölder continuous (see Lemma˜5.1), which implies that the term $\int_{0}^{t}c(s)\mathrm{d}H_{s}$ can be defined in the sense of the Young integral. Therefore, we can claim that (2.8) is well defined, given that (2.9) is a linear equation. In this way, we also show that $n(X^{n}-X)$ indeed converges to a nontrivial limit and thus that order $n$ is optimal.
    Let us also point out that the general idea of this part is close to that in [7]. However, here we are dealing with additive fBM noise, whereas [7] considers central limit theorem type results that heavily rely on the multiplicative structure of the noise; therefore, the detailed and essential analysis deviates. In the end, unlike the weakly convergence result from [7], we are able to show that $n(X^{n}_{t}-X_{t})$ converges to $c(t)$ in probability.